OpenStudy (anonymous):

Solve for x. 2/x = 6/2 2/3 1 1/2 6

OpenStudy (anonymous):

@marshallinwashington

OpenStudy (studygurl14):

cross multiply \(\large\frac{2}{x}=\frac{6}{2}\)

OpenStudy (studygurl14):

\(2(2)=6x\) \(4=6x\) \(\large\frac{4}{6}=x\) \(\large\frac{2}{3}=x\)

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

so its 2/3?

OpenStudy (anonymous):

Or .66666666666 forever. 2/3 is the exact answer. Point 666666 forever is an approximate answer.

OpenStudy (studygurl14):

yea

OpenStudy (anonymous):

k thnx:))